OpenStudy (anonymous):

which point lies on the line described by the equation below? y+11=-14(x-2)

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A. (11,-2) B. (2,-11) C.(-11,2) D. (-2,11)

OpenStudy (anonymous):

Well once we put the equation into slope-intercept it becomes y=-14x+17. We can put put them one by one into the equation, and if they are correct, it would equal out. The only one that does that is B. -11=-14(2)+17; -11=-28+17; -11=-11 And yeah that's one way of figuring it out.